Eddie is 7 years older than Brian. If Brian is x years old, then how old was Eddie 11 years ago?
(a) x - 18 (b) x - 4 (c) x - 7 (d) 7x - 11 (e) x + 18

Answer by ankor@dixie-net.com(22740)

Eddie is 7 years older than Brian.
If Brian is x years old, then how old was Eddie 11 years ago?
:
e = x + 7
:
e = (x+7) - 11
:
e = (x - 4), 11 yrs ago